OpenStudy (anonymous):

If you have 120 meters of fencing and want to enclose a rectangular area up against a long, straight wall, what is the largest area you can enclose?

OpenStudy (zzr0ck3r):

\(2x+2y=120\) and \(f(x) = xy\implies f(x)=x(60-x)\implies f '(x)=60-2x\)

OpenStudy (zzr0ck3r):

with me?

OpenStudy (anonymous):

yea I think so

OpenStudy (zzr0ck3r):

perimeter is given by p=2x+2y and they told us we have 120 to use so 120=2x+2y this is the same as 60=x+y which implies y = 60-x now area is given by A = xy but y = 60-x so we sub n for y A=x(60-x) this is what we want to maximize, so we take the derivative and get f'(x) = 60-2x right?

OpenStudy (zzr0ck3r):

that should say sun in not "sub n"

OpenStudy (zzr0ck3r):

ok now we set the derivative equal to 0 60-2x=0 which gives x=30 so we need to find y, so we use 2x+2y=120 which is again the same as x+y = 60 if x is 30 we know y is also then 30 , so our max area is 30*30=900

OpenStudy (anonymous):

Thank you so much, I just could not figure out how to set it up. that was very helpful!

OpenStudy (anonymous):

Oh my gosh, so if you are using 120 meters and we found the largest area to enclose what are the units for that?

OpenStudy (anonymous):

m^2?

OpenStudy (mathmale):

Yes, the units of measurement of area in this case would be m^2 (meters squared).